# ChineseSimpleVQA - “See the World, Discover Knowledge”: A Chinese Factuality Evaluation for Large Vision Language Models

### v3

@@ Abstract
The evaluation of factual accuracy in large vision language models (LVLMs) has lagged behind their rapid development, making it challenging to fully reflect these models’ knowledge capacity and reliability. In this paper, we introduce the first factuality-based visual question-answering benchmark in Chinese, named ChineseSimpleVQA, aimed at assessing the visual factuality of LVLMs across 8 major topics and 56 subtopics. The key features of this benchmark include a focus on the Chinese language, diverse knowledge types, a multi-hop question construction, high-quality data, static consistency, and easy-to-evaluate through short answers. Moreover, we contribute a rigorous data construction pipeline and decouple the visual factuality into two parts: seeing the world (i.e., object recognition) and discovering knowledge. This decoupling allows us to analyze the capability boundaries and execution mechanisms of LVLMs. Subsequently, we evaluate 35 advanced open-source and closed-source models, revealing critical performance gaps within this field.
ChineseSimpleVQA - “See the World, Discover Knowledge”:
@@ A Chinese Factuality Evaluation for Large Vision Language Models
Jihao Guthanks: Work done during an internship at Alibaba Group., Yingyao Wangthanks: Equal contribution., Pi Bu, Chen Wang, Ziming Wang, Tengtao Song, Donglai Wei, Jiale Yuan, Yingxiu Zhao, Yancheng He, Shilong Li, Jiaheng Liu, Meng Cao, Jun Songthanks: Corresponding Author., Yingshui Tan, Xiang Li, Wenbo Su, Zhicheng Zheng, Xiaoyong Zhu, Bo Zheng, Alibaba Group {gujihao.gjh, wangyingyao.wyy, jsong.sj}@taobao.com
@@ 2 Related Works
Several datasets also necessitate factual or commonsense knowledge Marino et al. (2019); Wang et al. (2015, 2017); Zellers et al. (2019). For instance, OK-VQA Marino et al. (2019) evaluates a model’s capability to answer questions using open-domain knowledge. Building upon this, A-OKVQA Schwenk et al. (2022) integrates diverse external knowledge and reasoning capabilities, while S3VQA Jain et al. (2021) introduces a dataset featuring questions that require object detection within images. Recently, factuality benchmarks have become increasingly important for LLMs: SimpleQA introduced by OpenAI Wei et al. (2024) assesses short-form factuality, and a Chinese SimpleQA He et al. (2024) is proposed by Alibaba Group.
@@ 3.4 Quality Control
Annotators must ensure question-answer pairs meet the standards by the following operations: 1) verifying and rewriting the questions or answers that do not meet the standards in Section3.3; 2) replacing unqualified or unrepresentative images with new ones from online sources; and 3) verifying answer accuracy via search engines. Pairs that cannot be improved are discarded. Finally, manual annotations are compared with LLM-verified results, retaining only fully consistent Q&A pairs. This rigorous process ensures accuracy and adherence to established standards.
@@ 4.2 Baseline Models
We evaluate 13 closed-source LVLMs (i.e., o1-preview (0901) , o1-mini (0901) , GPT-4o (0806) , GPT-4o-mini (0708)  [^4] https://openai.com/index/ , Gemini-1.5-pro-flash, Gemini-1.5-pro Team et al. (2024), Gemini-2.0-flash [^5] https://deepmind.google/technologies/gemini/flash/, Claude-3.5-sonnet2666https://www.anthropic.com/news/claude-3-5-sonnet, Claude-3.5-sonnet, Doubao-1.5-vision-pro [^7] https://team.doubao.com/en/direction/vision, Doubao-vision-pro/lite, Qwen-VL-max Bai et al. (2023)) and 22 open-source LVLMs (i.e., Qwen2-VL series Wang et al. (2024), Qwen2.5-VL series [^8] https://help.aliyun.com/zh/model-studio/developer-reference/use-qwen-by-calling-api,Deepseek-VL-2 series Wu et al. (2024), LLaVA-one-vision-0.6B/72B Li et al. (2024a), GLM-4V-9B [^9] https://open.bigmodel.cn/dev/api/normal-model/glm-4v, Llama-vision-11B/90B [^10] https://ollama.com/library/llama3.2-vision, InternVL series Chen et al. (2024).).
@@ 4.5 Performance on Different Topics
3We select the top 10 models with the highest CO rate on the final questions and evaluated their performance across different topics. The results are shown in Figure 3. For all topics, the trends in model performance for both recognition and final questions align with the conclusions presented in the previous section. It is obvious that these models are more adept at topics such as modern architecture, engineering, and science. However, their performance is slightly lower on natural topics, as these questions often need more specialized knowledge than common sense can provide.
@@ 7 Limitation
From the perspective of topic distribution, the coverage of topics discussed in this paper remains relatively limited. A broader topic coverage would contribute to a more comprehensive assessment of visual knowledge. In terms of data volume, after rigorous quality filtering and difficulty balancing, the final version of our dataset represents only a tiny fraction of the overall data. Therefore, expanding the dataset will be a key focus of our future efforts. Finally, regarding model structure, improving the visual factuality of the model to enhance its performance on ChineseSimpleVQA will become increasingly important. However, this issue may already be beyond the scope of this paper.
@@ B.3 Rankings on ChineseSimpleVQA
B.4 More Results

### v4

# “See the World, Discover Knowledge”: A Chinese Factuality Evaluation for Large Vision Language Models

@@ Abstract
The evaluation of factual accuracy in large vision language models (LVLMs) has lagged behind their rapid development, making it challenging to fully reflect these models’ knowledge capacity and reliability. In this paper, we introduce the first factuality-based visual question-answering benchmark in Chinese, named ChineseSimpleVQA, aimed at assessing the visual factuality of LVLMs across 8 major topics and 56 subtopics. The key features of this benchmark include a focus on the Chinese language, diverse knowledge types, a multi-hop question construction, high-quality data, static consistency, and easy-to-evaluate through short answers. Moreover, we contribute a rigorous data construction pipeline and decouple the visual factuality into two parts: seeing the world (i.e., object recognition) and discovering knowledge. This decoupling allows us to analyze the capability boundaries and execution mechanisms of LVLMs. Subsequently, we evaluate 35 advanced open-source and closed-source models, revealing critical performance gaps within this field. Our evaluation-friendly code  [^1] https://github.com/OpenStellarTeam/ChineseSimpleVQA and data  [^2] https://huggingface.co/datasets/OpenStellarTeam/Chinese-SimpleVQA have already been open-sourced.
“See the World, Discover Knowledge”:
@@ A Chinese Factuality Evaluation for Large Vision Language Models
Jihao Guthanks: Work done during an internship at Alibaba Group., Yingyao Wangthanks: Equal contribution., Pi Bu, Chen Wang, Ziming Wang, Tengtao Song, Donglai Wei, Jiale Yuan, Yingxiu Zhao, Yancheng He, Shilong Li, Jiaheng Liu, Meng Cao, Jun Songthanks: Corresponding Author., Yingshui Tan, Xiang Li, Wenbo Su, Zhicheng Zheng Xiaoyong Zhu, Bo Zheng, Alibaba Group {gujihao.gjh, wangyingyao.wyy, jsong.sj}@taobao.com
@@ 2 Related Works
Several datasets also necessitate factual or commonsense knowledge Marino et al. (2019); Wang et al. (2015, 2017); Zellers et al. (2019); tan2024chinese. For instance, OK-VQA Marino et al. (2019) evaluates a model’s capability to answer questions using open-domain knowledge. Building upon this, A-OKVQA Schwenk et al. (2022) integrates diverse external knowledge and reasoning capabilities, while S3VQA Jain et al. (2021) introduces a dataset featuring questions that require object detection within images. Recently, factuality benchmarks have become increasingly important for LLMs: SimpleQA introduced by OpenAI Wei et al. (2024) assesses short-form factuality, and a Chinese SimpleQA He et al. (2024) is proposed by Alibaba Group.
@@ 3.4 Quality Control
Annotators must ensure question-answer pairs meet the standards by the following operations: 1) verifying and rewriting the questions or answers that do not meet the standards in Section3.3; 2) replacing unqualified or unrepresentative images with new ones from online sources; and 3) verifying answer accuracy via search engines and Baidu Baike. Pairs that cannot be improved are discarded. Finally, manual annotations are compared with LLM-verified results, retaining only fully consistent Q&A pairs. This rigorous process ensures accuracy and adherence to established standards.
@@ 4.2 Baseline Models
We evaluate 13 closed-source LVLMs (i.e., o1-preview (0901) , o1-mini (0901) , GPT-4o (0806) , GPT-4o-mini (0708)  [^6] https://openai.com/index/ , Gemini-1.5-pro-flash, Gemini-1.5-pro Team et al. (2024), Gemini-2.0-flash [^7] https://deepmind.google/technologies/gemini/flash/, Claude-3.5-sonnet2888https://www.anthropic.com/news/claude-3-5-sonnet, Claude-3.5-sonnet, Doubao-1.5-vision-pro [^9] https://team.doubao.com/en/direction/vision, Doubao-vision-pro/lite, Qwen-VL-max Bai et al. (2023)) and 22 open-source LVLMs (i.e., Qwen2-VL series Wang et al. (2024), Qwen2.5-VL series [^10] https://help.aliyun.com/zh/model-studio/developer-reference/use-qwen-by-calling-api,Deepseek-VL-2 series Wu et al. (2024), LLaVA-one-vision-0.6B/72B Li et al. (2024a), GLM-4V-9B [^11] https://open.bigmodel.cn/dev/api/normal-model/glm-4v, Llama-vision-11B/90B [^12] https://ollama.com/library/llama3.2-vision, InternVL series Chen et al. (2024).).
@@ 4.5 Performance on Different Topics
We select the top 10 models with the highest CO rate on the final questions and evaluated their performance across different topics. The results are shown in Figure 3. For all topics, the trends in model performance for both recognition and final questions align with the conclusions presented in the previous section. It is obvious that these models are more adept at topics such as modern architecture, engineering, and science. However, their performance is slightly lower on natural topics, as these questions often need more specialized knowledge than common sense can provide.
@@ 7 Limitation
From the perspective of topic distribution, the coverage of topics discussed in this paper remains relatively limited. A broader topic coverage would contribute to a more comprehensive assessment of visual knowledge. In terms of data volume, after rigorous quality filtering and difficulty balancing, the final version of our dataset represents only a tiny fraction of the overall data. Therefore, expanding the dataset will be a key focus of our future efforts. Furthermore, as highlighted by this work jiang2025hiddendetect, datasets to evaluate the model against potential jailbreak attacks is crucial. Finally, regarding model structure, improving the visual factuality of the model to enhance its performance on ChineseSimpleVQA will become increasingly important. However, both performance improvement and security aspects may already be beyond the scope of this paper.
@@ Appendix A Overview of Appendix
  • Appendix B.4: Additional experimental analysis between popularity and difficulty.
  • Appendix B.5: Additional analysis of Human Calibration.
  • Appendix B.6: Additional experimental comparion with other Datasets.
  • Appendix B.7: Additional experimental comparison with Knowledge-Based Q&A
@@ B.3 Rankings on ChineseSimpleVQA
B.4 Analysis between Popularity and Difficulty
We select the first 100 samples from our dataset and monitor their page views on Wikipedia over the past 30 days to assess their popularity. We categorize the popularity into three ranges based on the number of views: (0-10), (10-50), and (50, +inf). We then evaluate the experimental performance of the following 10 closed-source models across these different levels of popularity as shown in Table 4. The results show that models generally perform better with higher popularity data, particularly in Recognition Q&A tasks. For data with over 50 views in the past 30 days, most models can achieve an accuracy rate of 100%.
B.5 Analysis of Human Calibration
We manually extract 110 samples of the main results from both the merged and recognition Q&A for verification and find that agreement with human assessments reached 95.5% and 97.3%, respectively. This proves the validity of using LLM-as-judge for evaluation.
B.6 Comparion with other Datasets
In order to explore whether models that perform well in other datasets continue to perform well in this dataset. We compare the evaluation performance of models on our ChineseSimpleVQA and text-based (no vision) short-form factuality benchmarks——ChineseSimpleQA He et al. (2024). The experimental results are shown in Table 5: As can be seen from the CO metric, the o1-preview model achieves the best results in the merged Q&A and Recognition Q&A of ChineseSimpleVQA, as well as in ChineseSimpleQA. In contrast, the gpt-4o-mini model performs the worst in the merged Q&A of ChineseSimpleVQA and in ChineseSimpleQA, demonstrating consistency. For the Recognition Q&A of ChineseSimpleVQA, the model that performed the worst was o1-mini. This indicates that our benchmark is a qualified factuality evaluation set that can also assess large vision language models’ image recognition capabilities. It holds broad evaluative applications for LVLMs.
B.7 Performance Comparison including Knowledge-Based Q&A
To assist us in analysis, we conduct an evaluation on text-only Knowledge-Based Q&A. The experimental results are shown in Table 6. From the CO metric perspective, we can find that the o1 preview still shows the best performance, which is consistent with the other two Q&A tasks. Additionally, the performance of Knowledge-Based Q&A is similar to Merged Q&A, and for most models, Knowledge-Based Q&A performs better. This indicates that for LVLMs, it is more challenging to simultaneously recognize images as well as answer Knowledge-Based questions, and compared to image recognition, Knowledge-Based Q&A is the more difficult part.
Model CO of Merged Q&A CO of Recognition Q&A
0 - 10 10 - 50 50 - +inf 0 - 10 10 - 50 50 - +inf
o1-preview 68.00 84.62 80.00 84.00 90.38 100.00
o1-mini 68.00 73.08 60.00 72.00 76.92 80.00
GPT-4o 72.00 73.08 80.00 76.00 86.54 100.00
GPT-4o-mini 48.00 69.23 80.00 68.00 92.31 80.00
Gemini-2.0-flash 72.00 78.85 60.00 84.00 90.38 80.00
Gemini-1.5-pro-flash 80.00 71.15 70.00 88.00 78.85 80.00
Gemini-1.5-pro 80.00 76.92 100.00 92.00 86.54 100.00
Claude-3.5-sonnet2 84.00 73.08 80.00 88.00 84.62 100.00
Claude-3.5-sonnet 68.00 67.31 80.00 92.00 78.85 90.00
Qwen-VL-max 88.00 73.08 90.00 80.00 78.85 90.00
Table 4: Different popularity’s difficulty of closed source LVLMs on multi-hop QAs (i.e. Merged Q&A and Recognition Q&A).
Model Overall results of Merged Q&A Overall results of Recognition Q&A Overall results of ChineseSimpleQA
[…]
o1-preview 68.8 24.6 6.5 73.6 71.1 79.1 13.6 7.3 85.3 82.1 63.8 24.0 12.2 72.7 67.9
o1-mini 52.7 38.1 9.2 58.1 55.3 64.8 24.9 10.3 72.2 68.3 39.5 39.9 20.6 49.7 44.1
GPT-4o 59.1 35.5 5.4 62.4 60.7 77.5 15.5 7.0 83.4 80.4 59.3 39.3 1.4 60.1 59.7
GPT-4o-mini 51.0 43.6 5.4 53.9 52.4 70.8 23.1 6.1 75.4 73.0 37.6 61.5 0.9 37.9 37.8
Gemini-1.5-pro 66.2 31.4 2.5 67.8 67.0 77.5 20.7 1.7 78.9 78.2 37.4 37.6 8.0 59.1 56.7
Claude-3.5-sonnet 59.5 26.4 14.2 69.4 64.0 69.5 20.2 10.3 77.5 73.3 45.2 26.4 27.4 63.6 53.5
Table 5: Performance comparison between ChineseSimpleVQA and ChinsesSimpleQA He et al. (2024). For metrics, CO, NA, IN, and CGA denote “Correct”, “Not attempted”, “Incorrect”, and “Correct given attempted”, respectively. The highest scores among models in each metric are highlighted in green.
Model Overall results of Merged Q&A Overall results of Recognition Q&A Overall results of Knowledge-Based Q&A
[…]
o1-preview 68.8 24.6 6.5 73.6 71.1 79.1 13.6 7.3 85.3 82.1 72.9 21.0 6.6 77.5 74.8
o1-mini 52.7 38.1 9.2 58.1 55.3 64.8 24.9 10.3 72.2 68.3 50.9 34.2 14.9 59.8 55.0
GPT-4o 59.1 35.5 5.4 62.4 60.7 77.5 15.5 7.0 83.4 80.4 66.5 31.6 1.8 67.8 67.2
GPT-4o-mini 51.0 43.6 5.4 53.9 52.4 70.8 23.1 6.1 75.4 73.0 47.6 51.5 1.2 47.9 47.2
Gemini-2.0-flash 64.5 29.5 5.9 68.6 66.5 76.7 19.6 3.7 79.7 78.2 69.4 30.3 0.5 69.6 69.4
Gemini-1.5-pro-flash 56.5 34.6 8.8 62.0 59.2 70.3 25.9 3.8 73.1 71.6 57.4 38.5 4.4 59.8 58.5
Gemini-1.5-pro 66.2 31.4 2.5 67.8 67.0 77.5 20.7 1.7 78.9 78.2 66.5 32.5 1.5 67.0 66.5
Claude-3.5-sonnet2 63.8 30.6 5.5 67.6 65.6 77.6 17.2 5.2 81.9 79.7 62.4 30.0 7.9 67.4 64.6
Claude-3.5-sonnet 59.5 26.4 14.2 69.4 64.0 69.5 20.2 10.3 77.5 73.3 56.5 26.7 16.7 67.9 61.7
Qwen-VL-max 56.5 39.6 3.8 58.8 57.6 72.9 24.6 2.5 74.7 73.8 58.0 39.5 1.7 59.9 59.3
Table 6: Performance comparison of closed source and open source LVLMs on multi-hop QAs (i.e. Merged Q&A, Knowledge-Based Q&A and Recognition Q&A). For metrics, CO, NA, IN, and CGA denote “Correct”, “Not attempted”, “Incorrect”, and “Correct given attempted”, respectively. The highest scores among models in each metric are highlighted in green.
Figure 12: Rankings of different models on ChineseSimpleVQA.
B.8 More Results